OpenStudy (anonymous):

Write each function in vertex form. Y=x^2+3x

OpenStudy (anonymous):

\[ Y=x^2+3x + \frac 9 4 - \frac 9 4\\ = \left ( x + \frac 3 2\right)^2- \frac 94 \]